# Sard properties for polynomial maps in infinite dimension

### v2

@@ 1.3. Applications to Endpoint maps of Carnot groups
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Given $\ell\in\mathbb{N}$, there exists $r=r(m,s,k,\ell)>0$ such that, letting $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, with $I=[0,1]$, it holds
(41) $\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap V)\big{)}=\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{V})\big{)}=0.$
Namely, the Sard property holds on the space of piecewise real–analytic controls with radius of convergence $>r$ and with $\ell$ pieces. Furthermore, letting $W=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell)$, then it holds
(42) $\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap W)\big{)}\leq\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{W})\big{)}\leq m-1.$
@@
Sussmann established that for real–analytic sub-Riemannian structures (such as Carnot groups), minimizing horizontal paths are real–analytic on an open and dense set of their interval of definition [Sus14]. In light of that result, Theorem J hints at a unexpected link between the regularity problem of sub-Riemannian geodesics and the minimizing Sard conjecture.
@@
Of course a subset of controls $V\subset L^{2}(I,\mathbb{R}^{k})$ must satisfy $\mathcal{E}(V)=\mathbb{G}$ to be relevant for the theory. By the Rashevskii-Chow theorem, $\mathcal{E}:L^{2}(I,\mathbb{R}^{k})\to\mathbb{R}^{m}$ is surjective. We observe that $\mathcal{E}$ is surjective also on $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ if $\ell\geq 4m$, since from the proof of the Rashevskii-Chow theorem any pair of points can be joined by a concatenation of $4m$ horizontal paths with constant controls, see [ABB20, Sec. 3.2]. Actually, the Endpoint map is surjective when restricted on any set that is dense in $L^{2}(I,\mathbb{R}^{k})$, see e.g. [Bel96, Sec. 2.5]. It follows that $\mathcal{E}$ is surjective on the sets of Theorem J since in any case they contain all (restrictions on $I$ of) polynomial controls.
Finally, we prove in Theorem 52 that the Endpoint map on Carnot groups is surjective already when restricted to polynomial controls of large enough degree (depending on $\mathbb{G}$). This is a particular case of the following more general result (see Theorem 53).
Theorem K (Quantitative surjectivity).
@@
In other words, one can find an $(m+1)$-dimensional vector subspace $E\subset L^{2}(I,\mathbb{R}^{k})$ of controls (depending on $\mathbb{G}$) such that the restriction of the Endpoint map to $E$ is surjective. Furthermore, $E$ can be assumed to have generators in any prescribed dense set $S\subset L^{2}(I,\mathbb{R}^{k})$. Theorem K can be seen as a more refined version (for Carnot groups) of the well–known fact that the Endpoint map is surjective when restricted on any set of controls that is dense in $L^{2}(I,\mathbb{R}^{k})$, see e.g. [Bel96, Sec. 2.5].
@@ 1.4. Closing thoughts on the Sard conjecture
We do not venture in a guess in favor of Sard conjecture on the whole $H=L^{2}(I,\mathbb{R}^{k})$, instead we propose a direction of future investigation that can be approached with our techniques. Recall that $H_{\mathrm{min}}\subset H$ is the set of energy-minimizing controls. It is known that $H_{\min}$ is boundedly compact in $H$ i.e., the intersection $H_{\min}\cap B_{H}(r)$ is compact for all $r>0$, see [ABB20, Thm. 8.66], [Agr98]. In fact, from the very recent result by Lokutsievskiy and Zelikin [LZ23, Cor. 1 $(\mathfrak{G})$], one can deduce the following estimate for the $n$–width on Carnot groups
@@
for some $C>0$ depending on $\mathbb{G}$. Estimate (44) can be understood as a quantitative compactness property for $H_{\min}$. Unfortunately, the polynomial decay (w.r.t. $n$) of (44) is too weak to apply Theorem A, since the latter requires an exponential one.
@@
then Theorem F would settle the minimizing Sard conjecture on Carnot groups.
We also note that it would be sufficient to prove (45) for the set of the so-called strictly abnormal energy-minimizing controls $H_{\min}^{\mathrm{str.abn}}\subsetneq H_{\min}$. Unfortunately, we were not able to produce a direct estimate of the $n$–width of (bounded subsets of) $H_{\min}$ or $H_{\min}^{\mathrm{str.abn}}$, so that we have no further evidence to support this idea towards a proof of the minimizing Sard conjecture.
@@ 2. Table of notations
{xltabular}
p0.19 X +$\mathbb{R}^{m}_{+}$ $m$-tuples of positive real numbers
$B_{X}(x,r)$ closed ball of radius $r>0$ centered at $x$ of a Banach space $X$
$B_{X}(r)$ closed ball of radius $r>0$ centered at 00 of a Banach space $X$
$B_{X}$ closed unit ball centered at 00 of a Banach space $X$
$\mathcal{L}(X,Y)$ linear continuous maps between banach spaces $X$, $Y$
$\|\cdot\|_{\operatorname{op}}$ operator norm on $\mathcal{L}(X,Y)$, see (119)
$\|\cdot\|_{X}$ norm for a Banach space $X$ (the $X$ is omitted when there is no confusion)
$\mathcal{B}(X,Y)$ continuous bounded maps between Banach spaces $X$, $Y$
$\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ piece-wise real–analytic functions, Section 5.2
$\mu$ Lebesgue measure
$\Omega_{n}(K,H)$ Kolmogorov $n$–width of $K\subset H$, Definition 26
$\omega(K,H)$ asymptotic Kolmogorov width of $K\subset H$, Eq. 25
$\mathcal{P}_{d}^{m}(H)$ set of well-approximable maps, Definition 1
$\beta_{0}(d,m)$ semialgebraic constant, Remark 2
$\omega(d,m)$ threshold constant, Theorem F
$\mathcal{H}^{i}(A)$ $i$-dimensional Hausdorff measure of a rectifiable Borel set $A$
$V_{i}(A)$ Vitushkin variation of a bounded semialgebraic set $A$, Definition 14
$b_{0}(S)$ number of connected components of a set $S$
$\sigma_{k}(L)$ $k$-th singular value of a linear map $L:H\to\mathbb{R}^{m}$ on a Hilbert space, ordered in decreasing order $\sigma_{1}(M)\geq\dots\geq\sigma_{m}$, Definition 16
$\operatorname{Crit}(f)$ critical points of a map, Eq. 115
$\operatorname{Crit}_{\nu}(f)$ $\nu$-critical points of a map, Eq. 115
$C^{\Lambda}(f)$ $\Lambda$-critical (or almost-critical) points of a map, Definition 19
$M(\varepsilon,S)$ $\varepsilon$-entropy of a set $S\subset\mathbb{R}^{m}$, Definition 24
$\dim_{e}(S)$ entropy dimension of a set $S\subset\mathbb{R}^{m}$, Definition 24
$\mathcal{U}_{\varepsilon}(S)$ $\varepsilon$-neighbourhood of a subset $S\subset\mathbb{R}^{n}$
$\mathrm{dist}_{H}$ Hausdorff distance
$\beta_{\mathrm{tm}}$ Thom-Milnor bound constant, Theorem 12
$\beta_{\mathrm{dc}}$ definable choice constant, Theorem 13
$\beta_{\mathrm{yc}}$ Yomdin-Comte constant, Theorem 18
$D(S)$ diagram of a semialgebraic set $S$, Definition 10
$\mathbb{G}$ Carnot group (of step $s$, rank $k$, dimension $m$), Definition 43
$\mathcal{E}$ Endpoint map, Definition 45
@@ 3.1. Semialgebraic sets and maps
The goal of this section is to prove Theorem G. We begin by recalling some basic notions from semialgebraic geometry.
@@
is a diagram for $S$. Below, the equation “$D(S)=(m,c,d)$” will mean that there exists a representation of $S$ as in (47) with $n\leq m$, $a\cdot\max_{i}\{{b_{i}}\}\leq c$ and $\max_{i,j}\{\deg(p_{ij})\}\leq d$.
@@ 3.1.1. Dimension and stratifications
Every semialgebraic subset of $\mathbb{R}^{n}$ can be written as a finite union of semialgebraic sets, each of them semialgebraically homeomorphic to an open cube $(0,1)^{m}\subset\mathbb{R}^{m}$, for some $m\leq n$ ([BCR98, Thm. 2.3.6]). This allows to define the dimension of a semialgebraic set as the maximum of the dimensions $m$ of these cubes. Every semialgebraic set can be written as a finite union of smooth, semialgebraic disjoint manifolds called strata ([BCR98, Prop. 9.1.8]).
@@ 3.1.2. Semialgebraic triviality
This result is called semialgebraic triviality, see [BPR06, Thm. 5.46].
@@ 3.1.3. Projections
The image of a semialgebraic set $S\subset\mathbb{R}^{n}\times\mathbb{R}^{\ell}$ under the projection map $\pi:\mathbb{R}^{n}\times\mathbb{R}^{\ell}\to\mathbb{R}^{n}$ is semialgebraic ([BCR98, Thm. 2.2.1]). Moreover, it follows from [BPR06, Thm. 14.16, Notation 8.2] [^1] Note that the role of $k$ and $\ell$ for us is swapped with respect to [BPR06, Thm. 14.16]: we eliminate $\ell$ variables from a list of $n+\ell$, they eliminate $k$ variables from a list of $k+\ell$. that there exists $C_{1}>0$ such that if $(n+\ell,c,d)$ is a diagram for $S$, we can write
@@ 3.1.4. Thom-Milnor bound
We will need the following qualitative bound on the number of connected components of a semialgebraic set, that follows from [BPR06, Thm. 7.50].
@@ 3.2. Definable choice
We recall from [LRT24] the following quantitative approximate definable choice result, which will be a key tool in the proof of Theorem G.
Let us explain the name and the context of this result. Given a polynomial map $F:\mathbb{R}^{n}\to\mathbb{R}^{\ell}$ and a compact semialgebraic set $A\subset\mathbb{R}^{n}$, for every $y\in F(A)$ one would like to choose a point $x(y)\in F^{-1}(y)\cap A$ so that the resulting set $C:=\{x(y)\}_{y\in F(A)}$ is semialgebraic and with controlled dimension. In particular, the “choice” of such $x(y)$ has to be “definable”. In this way $C$ would be of dimension $\ell$. The fact that this can be done follows from semialgebraic triviality (see Section 3.1.2). However, in this way there is no control on the geometry of $C$ (in particular, the number of its connected components) in terms of the data defining $A$ (namely, its diagram, see Definition 10).
@@
Theorem 13 (Quantitative approximate definable choice [LRT24]).
@@ 3.3. Variations and their behaviour under polynomial maps
We recall the definition of variations, introduced by Vitushkin [Vit55, Vit61] and developed in the semialgebraic context by Comte and Yomdin [YC04]. The variations encode the size of a set, through an integral-geometric approach. Given a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ and a semialgebraic set $A\subset\mathbb{R}^{n}$, it is possible to estimate the variations of the set $p(A)$, in a quantitative way, see [YC04, Thm. 7.2]. However, these estimates are not quantitative with respect to the dimensional parameter $n$. The main result of this section is Theorem 18, where we prove a version of [YC04, Thm. 7.2], where the dependence on $n$ is explicit, thanks to the use of Theorem 13.
@@
Let $A\subset\mathbb{R}^{n}$ be a bounded semialgebraic set. We define the 00–th variation of $A$ as
@@
where $b_{0}(A)$ denotes the number of connected components of $A$. For $i=1,\ldots,n$ we define the $i$–th variation of $A$ as
@@
We note that the integrand in (64) is measurable, see [YC04, Sec. 3].
A list of properties of variations is given in [YC04, page 35]. Here we prove the following one.
@@
We recall a classical result in Integral Geometry, called Cauchy–Crofton formula [Fed47, Thm. 5.14]: if $A$ is a $i$-rectifiable and Borel subset of $\mathbb{R}^{n}$, for $i\leq n$, then it holds
@@
If $A\subset\mathbb{R}^{n}$ is semialgebraic, the set $A$ can be written as a finite union of smooth, semialgebraic disjoint manifolds called strata (see Section 3.1.1). We can relabel the strata so that
@@
where each $A_{j}$ is a finite union of smooth $j$-dimensional manifolds. For any fixed $Z\in G_{\ell}(\mathbb{R}^{N})$ and for almost every $x\in Z$ the set $\pi_{Z}^{-1}(x)$ does not intersect [^2] We note that in this step of the proof it is enough to use the Semialgebraic Sard’s theorem [BCR98, Thm. 9.6.2], which is logically independent from the result in the smooth category. $A_{j}$ for $j<\ell$. Hence, $V_{0}(A\cap\pi_{Z}^{-1}(x))=V_{0}(A_{\ell}\cap\pi_{Z}^{-1}(x))$ for almost every $x\in Z$. Furthermore, for almost every $x\in Z$ the intersection $A_{\ell}\cap\pi_{Z}^{-1}(x)$ is transverse, and $\dim A+\dim\pi_{Z}^{-1}(x)=n$, so that $A_{\ell}\cap\pi_{Z}^{-1}(x)$ has dimension zero. Thus, its cardinality coincides with its number of connected components. We conclude by (67). ∎
@@
In finite dimension, the following Weyl inequality holds for the singular values of the difference of two matrices (see [Tao12, Ex. 1.3.22 (iv)]): for $n\geq m$ and linear maps $L_{1},L_{2}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ we have
@@
In the next result, given a semialgebraic set $A\subset\mathbb{R}^{n}$ and a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$, we bound the $i$-th variation $V_{i}(p(A))$ in terms of the singular values of $p$ and the diagram of $A$. Its proof follows the blueprint of [YC04, Thm. 7.2]. The main novelty (which is key for our work) is that the dependence on the dimension $n$ is explicit.
@@
An inspection of the proof shows that it holds $\beta_{\mathrm{yc}}=\beta_{\mathrm{tm}}^{2}\beta_{\mathrm{dc}}$ where $\beta_{\mathrm{tm}}=\beta_{\mathrm{tm}}(c(m+1),d)$ is the constant from the Thom-Milnor-type bound of Theorem 12, and $\beta_{\mathrm{dc}}=\max_{i\leq m}\beta_{\mathrm{dc}}(c,d,i)$ is the definable choice constant from Theorem 13.
@@
In the course of the proof, see (99), we obtain the following estimate, reminiscent of an area formula for variations, which is an improved version of [YC04, Thm. 7.2], with the explicit dependence of the constants on $n$: there exist a constant $\beta_{\mathrm{yc}}=\beta_{\mathrm{yc}}(c,d,m)$, such that
@@
By Definition 14, for $i=1,\dots,m$, we have
@@
For any $x\in Z$, the set $A\cap(\pi_{Z}\circ p)^{-1}(x)$ is semialgebraic with controlled diagram (more precisely, $D(A\cap(\pi_{Z}\circ p)^{-1}(x))=(n,c(1+m),d)$. Using Theorem 12 we obtain
@@
where $\beta_{\mathrm{tm}}=\beta_{\mathrm{tm}}(c(1+m),d)$ depends only on $c,d,m$. Note that, since by definition $V_{0}=b_{0}$, (79) yields the thesis of the theorem for the case $i=0$. We continue then assuming $i=1,\dots,m$. Denoting by $\mathbbm{1}_{S}$ the characteristic function of a set $S$, we get

### v3

@@ 1.3. Applications to Endpoint maps of Carnot groups
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Let $I=[0,1]$. For $\ell\in\mathbb{N}$, there exists $r=r(m,s,k,\ell)>1$ such that
(42) $\mu\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell))\big)=\mu\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)})\big)=0.$
Namely, the Sard property holds on the space of piecewise real–analytic controls with radius of convergence >r>r and with $\ell$ pieces. Furthermore it holds
(43) $\dim_{\mathcal{H}}\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell))\big)\leq\dim_{\mathcal{H}}\big(\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell)})\big)\leq m-1.$
@@
Remark 7 (Stronger versions).
By inspecting the proof of Theorem˜J, one actually obtains more general estimates for the Hausdorff dimension of the sets of critical values as in Corollary˜I. Furthermore, all these results hold in a stronger form, replacing $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ with the larger space
(44) $\mathrm{span}\left(\overline{\{u\in\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)\mid\|u\|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)}\leq 1\}}\right),\qquad r\in(0,\infty],$
where $\|\cdot\|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)}$, defined in (280), is a sort of Hardy $\infty$-norm, while the closure is w.r.t. $L^{2}(I,\mathbb{R}^{k})$.
@@
Sussmann established that for real–analytic sub-Riemannian structures (such as Carnot groups), minimizing horizontal paths are real–analytic on an open and dense set of their interval of definition [Suss-realanal]. In light of that result, Theorem˜J hints at an unexpected link between the regularity problem of sub-Riemannian geodesics and the minimizing Sard conjecture.
@@
Of course a subset of controls $V\subset L^{2}(I,\mathbb{R}^{k})$ must satisfy $\mathcal{E}(V)=\mathbb{G}$ to be relevant for the theory. By the Rashevskii-Chow theorem, $\mathcal{E}:L^{2}(I,\mathbb{R}^{k})\to\mathbb{R}^{m}$ is surjective. We observe that $\mathcal{E}$ is surjective also on $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ if $\ell\geq 4m$, since from the proof of the Rashevskii-Chow theorem any pair of points can be joined by a concatenation of $4m$ horizontal paths with constant controls, see [ABB, Sec. 3.2]. Actually, the Endpoint map is surjective when restricted on any set that is dense in $L^{2}(I,\mathbb{R}^{k})$, see e.g. [Bellaiche, Sec. 2.5]. It follows that $\mathcal{E}$ is surjective on the sets of Theorem˜J since in any case they contain all (restrictions on $I$ of) polynomial controls.
Finally, we prove in Proposition˜58 that the Endpoint map on Carnot groups is surjective already when restricted to polynomial controls of large enough degree (depending on $\mathbb{G}$). This is a particular case of the following more general result (see Proposition˜59).
Proposition K (Quantitative surjectivity).
@@
In other words, one can find an $(m+1)$-dimensional vector subspace $E\subset L^{2}(I,\mathbb{R}^{k})$ of controls (depending on $\mathbb{G}$) such that the restriction of the Endpoint map to $E$ is surjective. Furthermore, $E$ can be assumed to have generators in any prescribed dense set $S\subset L^{2}(I,\mathbb{R}^{k})$. Proposition˜K can be seen as a more refined version (for Carnot groups) of the well–known fact that the Endpoint map is surjective when restricted on any set of controls that is dense in $L^{2}(I,\mathbb{R}^{k})$, see e.g. [Bellaiche, Sec. 2.5].
@@ 1.4. Closing thoughts on the Sard conjecture
We do not venture in a guess in favor of Sard conjecture on the whole $H=L^{2}(I,\mathbb{R}^{k})$, instead we propose a direction of future investigation that can be approached with our techniques. Recall that $H_{\mathrm{min}}\subset H$ is the set of energy-minimizing controls. It is known that $H_{\min}$ is boundedly compact in $H$ i.e., the intersection $H_{\min}\cap B_{H}(r)$ is compact for all $r>0$, see [ABB, Thm. 8.66], [AAA-Rendiconti]. In fact, from the very recent result by Lokutsievskiy and Zelikin [Lev, Cor. 1 $(\mathfrak{G})$], one can deduce the following estimate for the nn–width on Carnot groups
@@
for some $C>0$ depending on $\mathbb{G}$. Estimate (46) can be understood as a quantitative compactness property for $H_{\min}$. Unfortunately, the polynomial decay (w.r.t. $n$) of (46) is too weak to apply Theorem˜A, since the latter requires an exponential one.
@@
then Theorem˜F would settle the minimizing Sard conjecture on Carnot groups.
We also note that it would be sufficient to prove (47) for the set of the so-called strictly abnormal energy-minimizing controls $H_{\min}^{\mathrm{str.abn}}\subsetneq H_{\min}$. Unfortunately, we were not able to produce a direct estimate of the nn–width of (bounded subsets of) $H_{\min}$ or $H_{\min}^{\mathrm{str.abn}}$, so that we have no further evidence to support this idea towards a proof of the minimizing Sard conjecture.
@@ 2. Table of notations
$\mathbb{R}^{m}_{+}$ ...................... $m$-tuples of positive real numbers
$B_{X}(x,r)$ ...................... closed ball of radius $r>0$ centered at $x$ of a Banach space $X$
$B_{X}(r)$ ...................... closed ball of radius $r>0$ centered at 0 of a Banach space $X$
$B_{X}$ ...................... closed unit ball centered at 0 of a Banach space $X$
$\mathcal{L}(X,Y)$ ...................... linear continuous maps between banach spaces $X$, $Y$
$\|\cdot\|_{\operatorname{op}}$ ...................... operator norm on $\mathcal{L}(X,Y)$, see (121)
$\|\cdot\|_{X}$ ...................... norm for a Banach space $X$ (the $X$ is omitted when there is no confusion)
$\mathcal{B}(X,Y)$ ...................... continuous bounded maps between Banach spaces $X$, $Y$
$\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ ...................... piece-wise real–analytic functions, Section˜5.2
$\mu$ ...................... Lebesgue measure
$\Omega_{n}(K,H)$ ...................... Kolmogorov nn–width of $K\subset H$, Definition˜30
$\omega(K,H)$ ...................... asymptotic Kolmogorov width of $K\subset H$, Eq.˜26
$\mathcal{P}_{d}^{m}(H)$ ...................... set of well-approximable maps, Definition˜1
$\beta_{0}(d,m)$ ...................... semialgebraic constant, Remark˜2
$\omega(d,m)$ ...................... threshold constant, Theorem˜F
$\mathcal{H}^{i}(A)$ ...................... $i$-dimensional Hausdorff measure of a rectifiable Borel set AA
$V_{i}(A)$ ...................... Vitushkin variation of a bounded semialgebraic set $A$, Definition˜15
$b_{0}(S)$ ...................... number of connected components of a set $S$
$\sigma_{k}(L)$ ...................... $k$-th singular value of a linear map $L:H\to\mathbb{R}^{m}$ on a Hilbert space, ordered in decreasing order $\sigma_{1}(M)\geq\dots\geq\sigma_{m}$, Definition˜17
$\operatorname{Crit}(f)$ ...................... critical points of a map, Eq.˜117
$\operatorname{Crit}_{\nu}(f)$ ...................... $\nu$-critical points of a map, Eq.˜117
$C^{\Lambda}(f)$ ...................... $\Lambda$-critical (or almost-critical) points of a map, Definition˜22
$M(\varepsilon,S)$ ...................... $\varepsilon$-entropy of a set $S\subset\mathbb{R}^{m}$, Definition˜27
$\dim_{e}(S)$ ...................... entropy dimension of a set $S\subset\mathbb{R}^{m}$, Definition˜27
$\mathcal{U}_{\varepsilon}(S)$ ...................... $\varepsilon$-neighbourhood of a subset $S\subset\mathbb{R}^{n}$
$\mathrm{dist}_{H}$ ...................... Hausdorff distance
$\beta_{\mathrm{tm}}$ ...................... Thom-Milnor bound constant, Theorem˜13
$\beta_{\mathrm{dc}}$ ...................... definable choice constant, Theorem˜14
$\beta_{\mathrm{yc}}$ ...................... Yomdin-Comte constant, Theorem˜19
$D(S)$ ...................... diagram of a semialgebraic set $S$, Definition˜11
$\mathbb{G}$ ...................... Carnot group (of step $s$, rank $k$, dimension $m$), Definition˜48
$\mathcal{E}$ ...................... Endpoint map, Definition˜51
@@ 3.1. Semialgebraic sets and maps
The goal of this section is to prove Theorem˜G. We begin by recalling some basic notions from semialgebraic geometry.
@@
is a diagram for $S$. Below, the equation “D(S)=(m,c,d)D(S)=(m,c,d)” will mean that there exists a representation of $S$ as in (49) with $n\leq m$, $a\cdot\max_{i}\{{b_{i}}\}\leq c$ and $\max_{i,j}\{\deg(p_{ij})\}\leq d$.
@@ 3.1.1. Dimension and stratifications
Every semialgebraic subset of $\mathbb{R}^{n}$ can be written as a finite union of semialgebraic sets, each of them semialgebraically homeomorphic to an open cube $(0,1)^{m}\subset\mathbb{R}^{m}$, for some $m\leq n$ ([BCR, Thm. 2.3.6]). This allows to define the dimension of a semialgebraic set as the maximum of the dimensions $m$ of these cubes. Every semialgebraic set can be written as a finite union of smooth, semialgebraic disjoint manifolds called strata ([BCR, Prop. 9.1.8]).
@@ 3.1.2. Semialgebraic triviality
This result is called semialgebraic triviality, see [BasuPoRoyBook, Thm. 5.46].
@@ 3.1.3. Projections
The image of a semialgebraic set $S\subset\mathbb{R}^{n}\times\mathbb{R}^{\ell}$ under the projection map $\pi:\mathbb{R}^{n}\times\mathbb{R}^{\ell}\to\mathbb{R}^{n}$ is semialgebraic ([BCR, Thm. 2.2.1]). Moreover, it follows from [BasuPoRoyBook, Thm. 14.16, Notation 8.2] [^1] Note that the role of $k$ and $\ell$ for us is swapped with respect to [BasuPoRoyBook, Thm. 14.16]: we eliminate $\ell$ variables from a list of $n+\ell$, they eliminate $k$ variables from a list of $k+\ell$. that there exists $C_{1}>0$ such that if $(n+\ell,c,d)$ is a diagram for $S$, we can write
@@ 3.1.4. Thom-Milnor bound
We will need the following qualitative bound on the number of connected components of a semialgebraic set, that follows from [BasuPoRoyBook, Thm. 7.50].
@@ 3.2. Definable choice
We recall from [LRT-selection] the following quantitative approximate definable choice result, which will be a key tool in the proof of Theorem˜G.
Let us explain the name and the context of this result. Given a polynomial map $F:\mathbb{R}^{n}\to\mathbb{R}^{\ell}$ and a compact semialgebraic set $A\subset\mathbb{R}^{n}$, for every $y\in F(A)$ one would like to choose a point $x(y)\in F^{-1}(y)\cap A$ so that the resulting set $C:=\{x(y)\}_{y\in F(A)}$ is semialgebraic and with controlled dimension. In particular, the “choice” of such $x(y)$ has to be “definable”. In this way $C$ would be of dimension $\ell$. The fact that this can be done follows from semialgebraic triviality (see Section˜3.1.2). However, in this way there is no control on the geometry of $C$ (in particular, the number of its connected components) in terms of the data defining $A$ (namely, its diagram, see Definition˜11).
@@
Theorem 14 (Quantitative approximate definable choice [LRT-selection]).
@@ 3.3. Variations and their behaviour under polynomial maps
We recall the definition of variations, introduced by Vitushkin [Vit1, Vitbook] and developed in the semialgebraic context by Comte and Yomdin [ComteYomdin]. The variations encode the size of a set, through an integral-geometric approach. Given a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ and a semialgebraic set $A\subset\mathbb{R}^{n}$, it is possible to estimate the variations of the set $p(A)$, in a quantitative way, see [ComteYomdin, Thm. 7.2]. However, these estimates are not quantitative with respect to the dimensional parameter $n$. The main result of this section is Theorem˜19, where we prove a version of [ComteYomdin, Thm. 7.2], where the dependence on $n$ is explicit, thanks to the use of Theorem˜14.
@@
Let $A\subset\mathbb{R}^{n}$ be a bounded semialgebraic set. We define the 0–th variation of $A$ as
@@
where $b_{0}(A)$ denotes the number of connected components of $A$. For $i=1,\ldots,n$ we define the ii–th variation of $A$ as
@@
We note that the integrand in (66) is measurable, see [ComteYomdin, Sec. 3].
A list of properties of variations is given in [ComteYomdin, page 35]. Here we prove the following one.
@@
We recall a classical result in Integral Geometry, called Cauchy–Crofton formula [FedererTransAmerMathSoc, Thm. 5.14]: if $A$ is a $i$-rectifiable and Borel subset of $\mathbb{R}^{n}$, for $i\leq n$, then it holds
@@
If $A\subset\mathbb{R}^{n}$ is semialgebraic, the set $A$ can be written as a finite union of smooth, semialgebraic disjoint manifolds called strata (see Section˜3.1.1). We can relabel the strata so that
@@
where each $A_{j}$ is a finite union of smooth $j$-dimensional manifolds. For any fixed $Z\in G_{\ell}(\mathbb{R}^{N})$ and for almost every $x\in Z$ the set $\pi_{Z}^{-1}(x)$ does not intersect [^2] We note that in this step of the proof it is enough to use the Semialgebraic Sard’s theorem [BCR, Thm. 9.6.2], which is logically independent from the result in the smooth category. $A_{j}$ for $j<\ell$. Hence, $V_{0}(A\cap\pi_{Z}^{-1}(x))=V_{0}(A_{\ell}\cap\pi_{Z}^{-1}(x))$ for almost every $x\in Z$. Furthermore, for almost every $x\in Z$ the intersection $A_{\ell}\cap\pi_{Z}^{-1}(x)$ is transverse, and $\dim A+\dim\pi_{Z}^{-1}(x)=n$, so that $A_{\ell}\cap\pi_{Z}^{-1}(x)$ has dimension zero. Thus, its cardinality coincides with its number of connected components. We conclude by (69). ∎
@@
In finite dimension, the following Weyl inequality holds for the singular values of the difference of two matrices (see [TaoRMT2012, Ex. 1.3.22 (iv)]): for $n\geq m$ and linear maps $L_{1},L_{2}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ we have
@@
In the next result, given a semialgebraic set $A\subset\mathbb{R}^{n}$ and a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$, we bound the $i$-th variation $V_{i}(p(A))$ in terms of the singular values of $p$ and the diagram of $A$. Its proof follows the blueprint of [ComteYomdin, Thm. 7.2]. The main novelty (which is key for our work) is that the dependence on the dimension $n$ is explicit.
@@
An inspection of the proof shows that it holds $\beta_{\mathrm{yc}}=\beta_{\mathrm{tm}}^{2}\beta_{\mathrm{dc}}$ where $\beta_{\mathrm{tm}}=\beta_{\mathrm{tm}}(c(m+1),d)$ is the constant from the Thom-Milnor-type bound of Theorem˜13, and $\beta_{\mathrm{dc}}=\max_{i\leq m}\beta_{\mathrm{dc}}(c,d,i)$ is the definable choice constant from Theorem˜14.
@@
In the course of the proof, see (101), we obtain the following estimate, reminiscent of an area formula for variations, which is an improved version of [ComteYomdin, Thm. 7.2], with the explicit dependence of the constants on $n$: there exist a constant $\beta_{\mathrm{yc}}=\beta_{\mathrm{yc}}(c,d,m)$, such that
@@
By Definition˜15, $V_{0}=b_{0}$. Furthermore, $b_{0}(p(A))\leq b_{0}(A)$. Thus, Theorem˜13 yields the thesis of the theorem for the case $i=0$. By Definition˜15, for $i=1,\dots,m$, we have
@@
For any $x\in Z$, the set $A\cap(\pi_{Z}\circ p)^{-1}(x)$ is semialgebraic with controlled diagram (more precisely, $D(A\cap(\pi_{Z}\circ p)^{-1}(x))=(n,c(1+m),d)$. Using Theorem˜13 we obtain
@@
where $\beta_{\mathrm{tm}}=\beta_{\mathrm{tm}}(c(1+m),d)$ depends only on $c,d,m$. Denoting by $\mathbbm{1}_{S}$ the characteristic function of a set $S$, we get